OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 8) and a radius of 9. would my answer be: (x-0)^2 +(y-8)^2=6 ???

OpenStudy (anonymous):

Not 6.. I meant 9

OpenStudy (jdoe0001):

close \(\bf 3^2=9\qquad thus \\ \quad \\ (x-{\color{brown}{ 0}})^2+(y-{\color{blue}{ 8}})^2={\color{purple}{ 3}}^2 \qquad center\ ({\color{brown}{ 0}},{\color{blue}{8}})\qquad radius={\color{purple}{ 3}}\)

OpenStudy (anonymous):

Im sorry, just a bit confused. How did you get 3?

OpenStudy (jdoe0001):

ohh the radius is hmmm hold the mayo

OpenStudy (jdoe0001):

hmmm the radius is just 9 flat... I see thus \(\bf (x-{\color{brown}{ 0}})^2+(y-{\color{blue}{ 8}})^2={\color{purple}{ 9}}^2\) sorta misread above =(

OpenStudy (jdoe0001):

I was going to say, your "6" is correct, if you just put it upside-down =)

OpenStudy (anonymous):

its fine :) if i dont put the ^2 would it be considered wrong??

OpenStudy (anonymous):

on the radius

OpenStudy (jdoe0001):

yes the radius is there, but squared the wording above says that "the radius is 9" so that makes it \(\large = 9^2\)

OpenStudy (anonymous):

okay cool thanks for your help!

OpenStudy (jdoe0001):

yw